# DCFormer: Efficient 3D Vision-Language Modeling with Decomposed Convolution

### v1

@@ Abstract
Vision-language models (VLMs) align visual and textual representations, enabling high-performance zero-shot classification and image-text retrieval in 2D medical imaging. However, extending VLMs to 3D medical imaging remains computationally challenging. Existing 3D VLMs rely on Vision Transformers (ViTs), which are computationally expensive due to self-attention’s quadratic complexity, or 3D convolutions, which demand excessive parameters and FLOPs as kernel size increases. We introduce DCFormer, an efficient 3D medical image encoder that factorizes 3D convolutions into three parallel 1D convolutions along depth, height, and width. This design preserves spatial information while significantly reducing computational cost. Integrated into a CLIP-based vision-language framework, DCFormer is evaluated on CT-RATE, a dataset of 50,188 paired 3D chest CT volumes and radiology reports, for zero-shot multi-abnormality detection across 18 pathologies. Compared to ViT, ConvNeXt, PoolFormer, and TransUNet, DCFormer achieves superior efficiency and accuracy, with DCFormer-Tiny reaching 62.0% accuracy and a 46.3% F1-score while using significantly fewer parameters. These results highlight DCFormer’s potential for scalable, clinically deployable 3D medical VLMs. Our codes will be publicly available.
@@ Introduction
Deep learning has revolutionized medical imaging, enabling automated disease diagnosis, accurate prognosis, and personalized treatment planning [1, 2, 3, 4]. With the emergence of Convolutional Neural Networks (CNNs) [5, 6] and Vision Transformers (ViTs) [7], deep learning has achieved state-of-the-art performance in supervised tasks such as lesion detection and organ segmentation [8, 9, 10, 11, 12]. Despite these advancements, supervised models require training on large-scale manually annotated datasets and lack generalization across diverse medical imaging tasks and modalities.
@@
To mitigate the reliance on labeled data, self-supervised learning (SSL) methods such as masked autoencoders [13] have emerged as powerful techniques for extracting meaningful representations from unlabeled images [14, 15]. Despite their advantages, these methods still require supervised fine-tuning on small labeled datasets. Vision-language models (VLMs) such as CLIP (Contrastive Language-Image Pretraining) [16] offer a promising alternative by aligning visual and textual representations in a shared latent space, enabling zero-shot capabilities. Initially successful in natural image domains [17, 18], CLIP has since been adapted for various 2D medical imaging tasks, including zero-shot classification, prompt-driven segmentation, image-text retrieval, radiology report generation, and visual question answering [19, 20, 21, 22, 23, 24, 25].
@@
While CLIP-based approaches have advanced 2D medical imaging, their application to 3D imaging remains underexplored due to several key challenges. A major challenge is the scarcity of large-scale, open-source datasets containing 3D image volumes paired with textual reports. To address this, Hamamci et al. [26] introduced CT-RATE, a dataset consisting of 50,188 reconstructed 3D chest computed tomography (CT) volumes from 25,692 scans of 21,304 patients, along with their corresponding radiology reports. They also proposed CT-CLIP, a 3D VLM for chest CT that employs CT-ViT [27], a model incorporating a two-stage spatial and temporal transformer. While CT-CLIP performs well in zero-shot detection and image-text retrieval tasks, its use of self-attention mechanisms leads to significant computational overhead. Additionally, its image encoder follows a ViT-like patching strategy, directly downsampling images with a fixed patch size of 20. This strategy leads to feature loss, as it fails to extract hierarchical representations—an essential aspect of medical imaging. Prior studies have shown that hierarchical feature extraction is crucial for capturing spatial and contextual information in medical images [9, 10, 28]. Despite these limitations, CT-CLIP marks a significant step toward adapting foundation models for 3D medical imaging.
Another challenge is the computational complexity of adapting CLIP-like VLMs to 3D imaging modalities such as CT and magnetic resonance imaging (MRI). The high-resolution, volumetric nature of 3D medical data imposes substantial computational demands, particularly with standard ViT-based or 3D CNN-based encoders. While ViTs effectively capture global spatial relationships, their computational complexity scales quadratically with input size, making them inefficient for 3D volumes. Likewise, 3D convolutions, which are well-suited for capturing local spatial features, become prohibitively expensive with large kernel sizes ($k>3$). To mitigate this, depthwise convolutions [29] have been explored as an alternative, reducing both parameter count and computational cost. However, they can still be inefficient for 3D imaging when large kernels and high output channel dimensions are required. Figure 1 compares the parameters and FLOPs of 2D and 3D standard versus depthwise convolutions, showing that standard 3D convolutions demand significantly more resources than their 2D counterparts. Although depthwise convolutions improve efficiency, their computational cost scales cubically with kernel size, posing challenges for large 3D medical datasets where large receptive fields are needed to capture complex anatomical structures. Additionally, their cumulative use throughout the network significantly increases the overall computational burden.
@@
Figure 1: FLOP and parameter comparison of 2D and 3D convolutions. For simplicity, we set the number of output channels $C=32$ and depth dimension $D=32$
@@
To overcome these challenges of existing 3D VLMs, we propose DCFormer, a family of architectures designed for 3D medical imaging. DCFormer introduces decomposed convolutions as an alternative to self-attention and conventional 3D convolutions, significantly reducing computational complexity while preserving the ability to extract both local and global spatial features. By factorizing 3D convolutions into three 1D convolutions along each spatial dimension, DCFormer aims to balance efficiency and performance, significantly reducing model parameters and FLOPs (Figure 2). We integrate DCFormer into the CLIP framework, enabling efficient joint vision-language learning for 3D medical imaging. Our experiments demonstrate that DCFormer achieves competitive zero-shot performance on CT-RATE dataset while significantly reducing computational costs compared to existing state-of-the-art methods. We anticipate that DCFormer will serve as a stepping stone for advancing scalable and efficient image encoders for vision-language understanding and inspire further research into lightweight, high-performance architectures for clinical applications.
@@
Figure 2: FLOP and parameter comparison between 3D depthwise convolution and decomposed depthwise convolution.
@@ Results
We train a CLIP-based joint vision-language learning framework (see Developing a CLIP framework using DCFormer in the Methods section) using the CT-RATE dataset, consisting of 50,188 non-contrast 3D chest CT volumes paired with corresponding radiology text reports. For the text encoder, we use CXR-BERT [30]. For the image encoder, we comprehensively evaluate our proposed DCFormer architecture by benchmarking its zero-shot performance against several state-of-the-art models, including ViT [7], CT-ViT [27, 26], TransUNet [9], ConvNeXt [31], InceptionNeXt [32], and PoolFormer [33] (see SOTA image encoders in the Methods section).
@@
DCFormer nano 920.3K 34.21 60.4 41.9 27.2 62.8
convnext [31] nano 3.19M 31.92 62.2 39.4 26.7 55.1
poolformer [33] nano 2.79M 27.14 60.2 37.0 24.8 52.3
DCFormer naïve 5.85M 49.48 63.1 44.5 29.5 65.5
vit [7] naïve 11.10M 39.05 55 42.5 25.8 71.5
convnext [31] naive 15.63M 96.84 60.7 42.4 27.7 63.8
poolformer [33] naïve 11.31M 63.75 60.1 39.1 25.7 56.8
transunet [9] naïve 12.48M 118.9 58.6 41.4 26.5 56.0
DCFormer tiny 15.1M 168.2 62.0 46.3 29.7 70.1
vit [7] tiny 26.34M 86.43 61 43.2 28.0 64.8
convnext [31] tiny 31.59M 156.31 62.5 42.1 28.2 60.1
transunet [9] tiny 23.93M 207.5 61.5 35.8 24.7 48.7
poolformer [33] tiny 20.68M 117.46 61.8 38.3 26.0 53.5
ctvit [26] - 101.1M 160.5 62.9 44.3 29.3 65.7
Zero-shot multi-abnormality detection results
Once trained to maximize the similarity between image and text embeddings, the CLIP framework enables multi-abnormality detection by inputting each abnormality as a text prompt (see *Zero-shot multi-abnormality detection with DCFormer* in the Methods section). Specifically, we use the prompts "Abnormality is present." and "Abnormality is not present." for each of the 18 distinct abnormalities, following the ablation study by Hamamci et al. [26]. We then compute the normalized probability of each abnormality being present in the CT image and evaluate model performance using accuracy, F1 score, precision, and recall.
@@
Table 1 compares the zero-shot performance and computational efficiency of multiple models on the CT-RATE dataset for image volumes of size 512×512×256. DCFormer provides a family of models with configurations ranging from nano to tiny variants, with parameter counts between 920.3K and 15.1M and GFLOPs spanning 34.21G to 168.2G. As shown in Table 1, DCFormer consistently achieves superior accuracy, F1 score, precision, and recall across all configurations while maintaining computational efficiency.
For instance, the nano variant achieves higher F1 score, precision, and recall than ConvNeXt and PoolFormer, despite having fewer parameters and similar FLOPs. Similarly, the naïve variant outperforms ConvNeXt, PoolFormer, and TransUNet, demonstrating the efficiency and effectiveness of DCFormer’s architecture. Notably, DCFormer achieves these results with substantially fewer computational resources compared to models like TransUNet and CTViT, which require much larger parameter counts and FLOPs. For example, the naive variant of DCFormer delivers 63.1% accuracy, 44.5% F1 score, 29.5% precision, and 65.5% recall using just 5.85M parameters and 49.48 GFLOPs—far fewer parameters and similar FLOPs compared to CTViT’s 101.1M parameters and 160.5 GFLOPs and TransUnet’s 23.93M parameters and 207.5 GFLOPs.
This efficiency enables DCFormer to achieve robust performance with a simple yet effective architecture, demonstrating its scalability and suitability for applications requiring both high accuracy and computational efficiency. Overall, for zero-shot implementation, DCFormer outperforms other models in efficiency, with its lightweight variants achieving competitive results while significantly reducing computational overhead. DCFormer’s ability to outperform SOTA models with fewer parameters and FLOPs highlights its potential for applications requiring both high performance and computational efficiency.
@@ Discussion
In this paper, we introduce DCFormer, a novel vision encoder designed to efficiently process 3D medical images within a CLIP-based vision-language framework. DCFormer addresses key computational challenges in 3D CT analysis, where existing models struggle with high-resolution volumetric data due to the quadratic scaling of self-attention in vision transformers and the high computational cost of 3D convolutions. To mitigate these issues, DCFormer leverages decomposed 3D convolutions as a token mixer, reducing FLOPs and parameter count while preserving strong spatial feature extraction capabilities. This design enables superior performance in zero-shot predictions.
We evaluate DCFormer on CT-RATE, a large-scale dataset containing 50,188 reconstructed 3D chest CT volumes paired with radiology reports. The model is tested under the zero-shot setting, benchmarking its performance against state-of-the-art methods including ViT, ConvNeXt, PoolFormer, and TransUNet. Our primary objective is to develop a model with strong zero-shot capabilities for multi-abnormality detection, minimizing the need for task-specific fine-tuning. As shown in Table 1, DCFormer consistently outperforms competing models in accuracy, F1 score, precision, and recall, all while requiring significantly fewer parameters and computational resources. This efficiency is driven by decomposed convolutions, which factorize 3D convolutions into multiple 1D operations along each spatial dimension, striking an optimal balance between performance and efficiency. Unlike CT-ViT and other computationally intensive 3D models, DCFormer offers a lightweight and scalable solution for large-scale 3D medical imaging, making it particularly well-suited for clinical applications requiring real-time decision-making.
While DCFormer demonstrates promising results, several future directions could enhance its capabilities. First, advanced prompt engineering may further improve zero-shot performance by better aligning image-text embeddings for classification tasks. Second, incorporating larger and more diverse datasets across different medical imaging modalities—such as MRI and PET—could improve generalization and further validate DCFormer’s effectiveness. Finally, integrating DCFormer with other VLMs, such as LLaVA [34] and LISA [35], could expand its applications to visual question answering and prompt-driven 3D image segmentation.

### v2

@@ Abstract
Vision-language models (VLMs) have been widely applied to 2D medical image analysis due to their ability to align visual and textual representations. However, extending VLMs to 3D imaging remains computationally challenging. Existing 3D VLMs often rely on Vision Transformers (ViTs), which are computationally expensive due to the quadratic complexity of self-attention, or on 3D convolutions, which require large numbers of parameters and FLOPs as kernel size increases. We introduce DCFormer, an efficient 3D image encoder that factorizes 3D convolutions into three parallel 1D convolutions along the depth, height, and width dimensions. This design preserves spatial information while significantly reducing computational cost. Integrated into a CLIP-based vision-language framework, DCFormer is trained and evaluated on CT-RATE, a dataset of 50,188 paired 3D chest CT volumes and radiology reports. In zero-shot and fine-tuned detection of 18 pathologies, as well as in image–text retrieval tasks, DCFormer consistently outperforms state-of-the-art 3D vision encoders, including CT-ViT, ViT, ConvNeXt, PoolFormer, and TransUNet. These results highlight DCFormer’s potential for scalable, clinically deployable 3D medical VLMs. Our code is available at: https://github.com/mirthAI/DCFormer.
@@ Introduction
Deep learning has revolutionized medical imaging by enabling automated disease diagnosis, accurate prognosis, and personalized treatment planning [1, 2, 3, 4]. In particular, Convolutional Neural Networks (CNNs) [5, 6] and Vision Transformers (ViTs) [7] have driven state-of-the-art performance in many supervised tasks (e.g., lesion detection and organ segmentation [8, 9, 10, 11, 12]). However, these supervised methods require training on large-scale manually annotated datasets and often fail to generalize across different imaging modalities and tasks. This limitation has motivated the development of more scalable and generalizable learning paradigms.
@@
To reduce the dependency on extensive data labeling, self-supervised learning (SSL) methods such as masked autoencoders [13] have been developed to extract informative representations from unlabeled images [14, 15]. While effective, these SSL methods still typically require subsequent supervised fine-tuning on small labeled datasets. Vision-language models (VLMs) such as CLIP (Contrastive Language-Image Pretraining) [16] offer a promising alternative by aligning visual and textual representations in a shared latent space, enabling zero-shot capabilities. Initially developed for natural image domains [17, 18], CLIP has since been adapted for 2D medical imaging tasks, including zero-shot classification, prompt-based segmentation, image-text retrieval, radiology report generation, and visual question answering [19, 20, 21, 22, 23, 24, 25].
@@
Despite the success of VLMs in 2D, their application to 3D imaging remains underexplored due to several challenges. One major challenge is the scarcity of large-scale public datasets of 3D image volumes paired with text reports. Recently, Hamamci et al. [26] addressed this by releasing CT-RATE, a dataset consisting of 50,188 reconstructed 3D chest computed tomography (CT) volumes (from 25,692 scans of 21,304 patients) with corresponding radiology reports. They also introduced CT-CLIP, a 3D VLM for chest CT that uses a vision transformer encoder (CT-ViT [27]) with a two-stage spatial and temporal transformer architecture. Although CT-CLIP achieved promising results in zero-shot detection and image-text retrieval tasks, its reliance on self-attention for the image encoder leads to significant computational overhead. Additionally, CT-CLIP uses a ViT-like patch tokenization strategy that downsamples each volume using a fixed patch size of 20. This strategy causes a loss of fine-grained information, as it fails to extract hierarchical representations — an essential element in medical imaging. Prior work has shown that capturing multi-scale, hierarchical features is crucial for understanding complex spatial and contextual information in medical images [9, 10, 28]. Nonetheless, CT-CLIP marks a significant step toward bringing VLMs to 3D medical imaging.
@@
Figure 1: Parameter count and computational cost (FLOPs) comparison for 2D and 3D standard and depthwise convolutions across kernel sizes. For simplicity, the number of output channels is fixed at $C=32$, and the depth dimension is set to $D=32$.
Another barrier to 3D VLMs is the computational complexity of processing volumetric data (e.g., CT scans). High-resolution 3D image volumes impose heavy computational demands on standard ViT-based or 3D CNN-based encoders. While ViTs are effective at capturing global relationships, their computational complexity scales quadratically with input size, making them inefficient for 3D volumes. Similarly, 3D convolutions are well-suited for local feature extraction, but they can become prohibitively expensive as convolutional kernel size grows beyond 3. To mitigate this, depthwise convolutions [29] have been explored as an alternative, reducing both parameter count and computational cost. However, they can still be inefficient for 3D imaging when large kernels and high output channel dimensions are required. Figure 1 compares the parameters and FLOPs of 2D and 3D standard versus depthwise convolutions, showing that standard 3D convolutions demand significantly more resources than their 2D counterparts. Although depthwise convolutions improve efficiency, their computational cost scales cubically with kernel size, posing challenges for 3D medical imaging where large receptive fields are needed to capture complex anatomical structures. Additionally, stacking many such layers in a deep network compounds computational burden.
@@
To address these challenges of existing 3D VLMs, we propose DCFormer, a family of 3D vision-language encoder architectures for medical images. DCFormer replaces both self-attention and standard 3D convolutions with a decomposed convolution strategy that significantly reduces computational complexity while preserving the ability to extract both local and global spatial features. In essence, each 3D convolution in DCFormer is factorized into three 1D convolutions applied along the height, width, and depth dimensions (Figure 3). This factorization significantly reduces the number of parameters and FLOPs required (Figure 2), achieving a much better balance of efficiency and performance. We integrate DCFormer as the vision encoder into a CLIP framework to enable efficient image-text representation learning on 3D imaging data. In our experiments on the CT-RATE dataset, DCFormer achieves competitive or better performance in zero-shot pathology detection and image-text retrieval tasks, while significantly reducing computational costs compared to existing state-of-the-art methods. To our knowledge, DCFormer is the first 3D vision encoder to use decomposed convolutions to improve vision-language modeling. We anticipate that DCFormer can serve as a stepping stone toward scalable, efficient image encoders for vision-language understanding, and hope it will inspire further research into lightweight, high-performance architectures for clinical applications.
@@
Figure 2: Parameter count and computational cost (FLOPs) comparison for standard 3D depthwise convolution and decomposed depthwise convolution.
@@ Results
We trained a CLIP-based vision-language framework (see Methods: Developing a CLIP framework using DCFormer ) using the CT-RATE dataset, which consists of 50,188 non-contrast 3D chest CT volumes paired with corresponding free-text radiology reports. For the text encoder, we used a domain-specific BERT model (CXR-BERT [30]). For the image encoder, we evaluated different variants of our proposed DCFormer architecture (nano, naïve, tiny) against several state-of-the-art 3D vision encoders, including ViT3D [7], CT-ViT [27, 26], TransUNet [9], ConvNeXt [31], and PoolFormer [32]. We focused on three main tasks: zero-shot abnormality detection, fine-tuned abnormality detection, and image–text retrieval.
Zero-Shot Multi-Abnormality Detection
Once trained to align images and text in a shared latent space, our CLIP model can perform zero-shot detection of multiple pathologies in a chest CT volume by using each pathology name as a text prompt (Figure 5b)). We adopt an approach similar to CheXzero[21] and CT-CLIP[26]: for each of the 18 candidate abnormalities, we provide a positive (’{Pathology} is present.’) prompt and a negative (’{Pathology} is not present.’) prompt. We compute the cosine similarity between between the CT image embedding and each of the text prompt embeddings. These similarities are used to compute the normalized probability of each abnormality being present. We evaluate model performance using accuracy, F1 score, precision, and recall.
@@
DCFormer nano 0.92 34.21 60.4 41.9 27.2 62.8
naïve 5.85 49.48 63.1 44.5 29.5 65.5
tiny 15.1 168.2 62.0 46.3 29.7 70.1
ConvNeXt [31] nano 3.19 31.92 62.2 39.4 26.7 55.1
naïve 15.63 96.84 60.7 42.4 27.7 63.8
tiny 31.59 156.31 62.5 42.1 28.2 60.1
PoolFormer [32] nano 2.79 27.14 60.2 37.0 24.8 52.3
naïve 11.31 63.75 60.1 39.1 25.7 56.8
tiny 20.68 117.46 61.8 38.3 26.0 53.5
TransUNet [9] naïve 12.48 118.9 58.6 41.4 26.5 56.0
tiny 23.93 207.5 61.5 35.8 24.7 48.7
ViT [7] naïve 11.10 39.05 55.0 42.5 25.8 71.5
tiny 26.34 86.43 61.0 43.2 28.0 64.8
CTViT [26] - 101.1 160.5 62.9 44.3 29.3 65.7
@@
Table 1 compares the zero-shot performance and computational efficiency of multiple models on the CT-RATE dataset (image volume size 512×512×256). DCFormer provides a family of models with configurations ranging from nano to tiny variants. Across all configurations, the lightweight variants of DCFormer consistently achieve superior results while reducing computational overhead. For instance, the nano variant of DCFormer achieves higher F1 score, precision, and recall than the corresponding nano variants of ConvNeXt and PoolFormer. Similarly, the naïve variant of DCFormer achieves a higher F1 score than the naïve variants of ConvNeXt, PoolFormer, and TransUNet. Notably, DCFormer achieves these results with significantly improved computational efficiency. For example, the naïve variant of DCFormer delivers a 44.5% F1 score—outperforming CT-ViT as well as the tiny variants of ConvNeXt, PoolFormer, and TransUNet—while using substantially fewer parameters and FLOPs: 5.85 million parameters and 49.48 GFLOPs, compared to CT-ViT’s 101.1 million parameters and 160.5 GFLOPs. DCFormer’s ability to outperform state-of-the-art models with far fewer computational resources underscores its potential for applications demanding both high performance and efficiency.
Fine-Tuned Multi-Abnormality Detection
Table 2: Fine-tuning performance of models trained on CT-RATE dataset at the resolution of 512x512x256.
DCFormer nano 0.92 34.21 60.1 45.0 28.6 69.6
naïve 5.85 49.48 64.2 47.2 31.0 70.2
tiny 15.1 168.2 64.1 48.6 31.4 72.9
ConvNeXt [31] nano 3.19 31.92 56.3 43.5 26.2 70.1
naïve 15.63 96.84 59.0 44.2 27.2 67.6
tiny 31.59 156.31 60.4 44.8 28.2 68.2
PoolFormer [32] nano 2.79 27.14 52.6 41.2 24.1 67.6
naïve 11.31 63.75 53.9 41.9 24.8 68.7
tiny 20.68 117.46 54.2 42.2 25.0 68.8
TransUNet [9] naïve 12.48 118.9 60.2 44.6 29.4 70.1
tiny 23.93 207.5 50.2 39.4 22.7 66.1
ViT [7] naïve 11.10 39.05 56.2 43.1 26.1 69.8
tiny 26.34 86.43 59.4 44.9 28.0 70.8
CTViT [26] - 101.1 160.5 61.8 46.4 29.4 70.1
To further validate DCFormer  we fine-tuned each model by adding a single linear classification layer on top of the frozen image encoder, and trained on labeled data for the same 18 chest abnormalities (Figure 5c). We used a Binary Cross-Entropy (BCE) loss to allow independent prediction of each abnormality (i.e., treating it as a multi-label classification rather than multi-class). Table 2 summarizes the fine-tuning results on the test set. As expected, fine-tuning improves the F1 scores of all models compared to their zero-shot performance. Notably, DCFormer variants continue to outperform all other architectures when fine-tuned. The DCFormer-tiny model achieves the highest F1 score at 48.6%, surpassing the best competing model (CTViT, 46.4%). DCFormer-naïve also achieves a strong F1 score of 47.2%, outperforming all other fine-tuned models. These gains highlight DCFormer’s efficient and effective feature extraction when adapted to specific tasks.
Image-Text Retrieval
Finally, we evaluated DCFormer on bidirectional image-text retrieval tasks. We computed cosine similarity scores between each CT volume embedding and each report embedding, and evaluated retrieval performance using Recall@k. As shown in Table 3, DCFormer delivers strong performance in both image-to-text (IR) and text-to-image (TR) tasks. The DCFormer tiny variant achieves the highest Recall@50, with 11.31% for IR and 12.37% for TR. The naïve variant of DCFormer significantly outperforms the nano variant, highlighting the effective scaling capability of our model. These results demonstrate DCFormer’s ability to capture semantic relationships between visual and textual data. We attribute the strong retrieval results to DCFormer’s use of large-kernel decomposed convolutions, which provide broad spatial context while maintaining efficiency.
Table 3: Image-text retrieval performance of models trained on CT-RATE dataset at the resolution of 512x512x256.
Models Variant IR TR
R@1 R@10 R@50 R@1 R@10 R@50
DCFormer nano 0.0001 0.0135 0.0617 0.0001 0.0135 0.0707
naive 0.0026 0.0220 0.0954 0.0016 0.0214 0.0971
tiny 0.0053 0.0296 0.1131 0.0039 0.0280 0.1237
ConvNeXt[31] nano 0.0013 0.0102 0.0562 0.0029 0.0135 0.0599
naive 0.0016 0.0161 0.0780 0.0023 0.0214 0.0740
tiny 0.0026 0.0135 0.0694 0.0010 0.0171 0.0796
PoolFormer[32] nano 0.0016 0.0095 0.0480 0.0023 0.0164 0.0592
naive 0.0013 0.0134 0.0569 0.0026 0.0151 0.0635
tiny 0.0007 0.0171 0.0638 0.0016 0.0115 0.0648
TransUNet[9] naive 0.0023 0.0236 0.1000 0.0033 0.0269 0.1105
tiny 0.0016 0.0131 0.0589 0.0026 0.0151 0.0651
ViT[7] naive 0.0016 0.0112 0.0500 0.0007 0.0109 0.0563
tiny 0.0010 0.0167 0.0628 0.0013 0.0148 0.0737
CTViT[26] - 0.0046 0.0293 0.1096 0.0023 0.0260 0.0921
@@ Discussion
In this paper, we introduced DCFormer, a novel 3D vision encoder designed to efficiently process volumetric medical images within vision–language models. DCFormer addresses key computational challenges in 3D image analysis, where conventional models struggle due to the quadratic complexity of vision transformers and the computational burden of large-kernel 3D convolutions. By leveraging decomposed 3D convolutions as a token mixing mechanism, DCFormer drastically reduces parameter count while preserving strong spatial feature extraction capabilities.
We evaluated DCFormer on CT-RATE, a large-scale dataset comprising 50,188 3D chest CT volumes paired with corresponding radiology reports, benchmarking its performance against several state-of-the-art 3D vision encoders, including ViT, ConvNeXt, PoolFormer, TransUNet, and CT-ViT. Across tasks such as multi-abnormality detection (in both zero-shot and fine-tuned settings) and image–text retrieval, DCFormer consistently outperformed all competing models. This performance is directly attributable to its decomposed convolution architecture, which captures multi-scale spatial features without the computational burden of self-attention or full 3D convolutions. Unlike CT-ViT and other resource-intensive models, DCFormer achieves a favorable balance between accuracy and efficiency, making it a practical and scalable solution for volumetric medical imaging.
Importantly, DCFormer’s low computational cost and fast inference time make it well suited for clinical applications requiring real-time responsiveness—such as automated triage, abnormality flagging, or preliminary report generation. With only 5.85 million parameters, the DCFormer-naïve variant can run on standard hospital hardware without requiring specialized infrastructure, significantly lowering the barrier to deploying advanced AI in routine workflows. We envision DCFormer as a foundational component in next-generation clinical decision support systems, helping democratize access to multimodal AI and accelerating the adoption of vision–language tools in healthcare.
Beyond the CLIP framework, DCFormer has been independently validated as a versatile vision backbone in the recently proposed Med3DVLM architecture [33]. In that work, DCFormer was combined with SigLIP [34] and incorporated into the LLaVA [35] framework to enable instruction-tuned multimodal understanding over 3D medical volumes and free-text prompts. The resulting system demonstrated competitive or superior performance in 3D report generation, zero-shot retrieval, and visual question answering tasks. This broader validation highlights DCFormer’s generalizability across different vision-language learning paradigms and tasks, reinforcing its potential as a robust foundation for multimodal medical AI.
While DCFormer demonstrates strong potential, several avenues for future work remain. First, more sophisticated prompt engineering techniques could further improve zero-shot performance by better aligning the visual and textual embeddings for each clinical query. Second, incorporating additional and more diverse training data – including other imaging modalities like MRI and PET – could improve the model’s generalization and validate its effectiveness across different 3D imaging contexts. Finally, integrating DCFormer with emerging multimodal frameworks (such as LLaVA [35] for vision–language question answering or LISA [36] for segmentation) could broaden its application to tasks like interactive diagnosis and automated 3D image interpretation.